OpenStudy (anonymous):

I got B Solve 5^x + 5 = 9x (2 points) Answer x ≈ -3.63 x ≈ -0.28 x ≈ 0.07 x ≈ 13.69

OpenStudy (anonymous):

9^x

jimthompson5910 (jim_thompson5910):

is the equation \[\Large 5^{x+5} = 9^{x}\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok then this means.... \[\Large 5^{x+5} = 9^{x}\] \[\Large \ln\left(5^{x+5}\right) = \ln\left(9^{x}\right)\] \[\Large (x+5)\ln\left(5\right) = x*\ln\left(9\right)\] \[\Large x*\ln\left(5\right)+5*\ln\left(5\right) = x*\ln\left(9\right)\] \[\Large 5*\ln\left(5\right) = x*\ln\left(9\right)-x*\ln\left(5\right)\] \[\Large 5*\ln\left(5\right) = x\left(\ln\left(9\right)-\ln\left(5\right)\right)\] \[\Large \frac{5*\ln\left(5\right)}{\ln\left(9\right)-\ln\left(5\right)} = x\] \[\Large x = \frac{5*\ln\left(5\right)}{\ln\left(9\right)-\ln\left(5\right)}\] \[\Large x \approx 13.69066371\] \[\Large x \approx 13.69\]

OpenStudy (anonymous):

love you thanks

jimthompson5910 (jim_thompson5910):

glad to be of help